surface brightness σ ≡ \frac{energy}{sq.arcsec sec}
or \frac{photon}{sq.arcsec sec}
or \frac{magnitude}{sq.arcsec}
σ = \frac{counts from star}{area}

= \frac{1247 counts per sec}{9 sq. arcsec} = 138 counts per sec per sq. arcsec

E5 means 5 = 10 (\frac{a − b}{a})

Φ(L) dL = \frac{number of galaxies in [L, L + dL]}{V} Φ(M) dM = \frac{number of galaxies in [M, M + dM]}{V}

N_{total} = ∫ Φ(L) dL

= ∫ Φ(M) dM

Φ(L) = \frac{Φ^{∗}}{L^{∗}} (\frac{L}{L^{∗}} )^{α}e^{−L/L^{∗}}